OpenStudy (anonymous):

The degree measures if the angles if a triangle are consecutive even integers. What is the degree measure of each of the three angles?

OpenStudy (texaschic101):

consecutive even integers...there will be 3 because it is a triangle x, x + 2, x + 4 angles of a triangle add up to 180 degrees x + (x + 2) + (x + 4) = 180 -- combine like terms 3x + 6 = 180 3x = 180 - 6 3x = 174 x = 174/3 x = 58 x + 2 = 58 + 2 = 60 x + 4 = 58 + 4 = 62 check.. 58 + 60 + 62 = 180 180 = 180 (correct) so your 3 angles are : 58,60,62